### v3

# On action rate admissibility criteria

Abstract
We formulate new admissibility criteria for initial value problems motivated by the least action principle. These are applied to a two-dimensional Riemann initial value problem for the isentropic compressible Euler fluid flow. It is shown that the criterion prefers the 2-shock solution to solutions obtained by convex integration by Chiodaroli and Kreml or to the hybrid solutions recently constructed by Markfelder and Pellhammer.
MSC: 35L65 (primary), 35C06, 35D30, 35LQ35, 76N10 (secondary).
Keywords: Weak Solution; Riemann Problem; Least Action Principle; Convex Integration ; Admissible Solution; Entropy Rate.
1 Introduction
The paper [9] introduces a version of the least action principle as a selection criterion in initial value problems possessing non-unique solutions. Interest arises from examples that use convex integration techniques to obtain families of non-unique wild solutions to the Euler system (De Lellis and Székelyhidi [4, 5]), and non-unique solutions to the Navier-Stokes equations (Buckmaster and Vicol [1]). Convex integration is also employed by Chiodaroli and Kreml [2] to investigate the two-dimensional Riemann initial value problem for the isentropic compressible Euler system. They construct a large set of entropic wild solutions to the Riemann problem.
In the present paper we continue the discussion of admissibility criteria started in [9] in view of the recent results of Markfelder and Pellhammer [14]. Admissibility criteria in fluid dynamics include the principles of minimum potential energy or entropy, various rate of change criteria, and that of least action (LAAP) developed in [9]. The latter principle applied to the initial value problem considered in [2] selects the 2-shock solution provided the solutions obtained by convex integration are the only other ones admitted for comparison. On the other hand, Chiodaroli and Kreml [2] show that there exist Riemann data for which there are wild solutions that possess entropy rate lower than the 2-shock solution which therefore is not selected by Dafermos’s entropy rate criterion [3]. On the other hand, in particular cases, LAAP selects the 2-shock in preference to wild solutions.
LAAP, however, does not select the 2-shock solution when the set of comparison solutions is enlarged to include other weak solutions in addition to the 2-shock and convex integration solutions. Markfelder and Pellhammer [14] recently constructed hybrid solutions to the same Riemann initial value problem for which, subject to prescribed initial data, LAAP does not identify the 2-shock solutions as being uniquely admissible. In fact, the example indicates that LAAP does not provide an admissible solution to the Riemann problem. Consequently, when the 2-shock solution is preferred in this Riemann problem, an alternative to LAAP must be sought.
The main purpose of this paper therefore is to formulate a new admissibility criterion, called LAAP0, which for the Riemann problem considered in [2] selects the 2-shock solution in comparison to the convex integration and hybrid solutions. LAAP0, while still motivated by the least action principle, involves a rate of change of action. Both the Dafermos entropy rate criterion and LAAP0 are local in time. The solutions constructed in [14] are excluded as they rely upon the global nature of LAAP.
Section 2 collects basic definitions, states the least action admissibility principle (LAAP) derived in [9] and formulates the new least action admissibility principle (LAAP0) and two closely related rate criteria. Section 3, after a brief description of convex integration, introduces the two-dimensional Riemann initial value problem for the barotropic and isentropic compressible Euler equations of gas dynamics and summarises further material from [2] including the construction of sub-solutions. The 2-shock solution is also described. Justification for the introduction of LAAP0 is provided in Section 4, while Section 5 defines the relevant notion of action and applies LAAP0 to the Riemann problem treated in [2] to show that the 2-shock solution is admissible.
2 Least action admissibility criteria
Let $S$ be a set of solutions ${\boldsymbol{u}}$ of the Cauchy problem for the evolutionary system specified by
$\frac{d{\boldsymbol{u}}}{dt}=F({\boldsymbol{u}}),\qquad{\boldsymbol{u}}(t_{0})={\boldsymbol{u}}_{0},\qquad t\in[t_{0},\,T],$ (2.1)
where $F({\boldsymbol{u}})$ is a given operator and $[t_{0},\,T]$ is the maximal time interval of existence of solutions belonging to $S$. Global existence is not assumed so that possibly $T<\infty.$ The system (2.1) is taken to be the Euler-Lagrange equation derived from a least action principle. Suppose that $S$ is not a singleton and in consequence (2.1) has non-unique solutions. The objective is to formulate an admissibility criterion defined in terms of the action that selects (not necessarily unique) elements of $S$. The specific form of action we will use is postponed to Section 5. For present purposes, it is sufficient to assume that for each solution ${\boldsymbol{u}}\in S$, the action $A({\boldsymbol{u}})(t_{0},\,t)$ is a non-decreasing function of time $t$ that vanishes initially for all ${\boldsymbol{u}}\in S$.
Further definitions are required, but before their statement, it is convenient to recall the Least Action Admissibility Principle (LAAP) introduced in [9] and applied to the Riemann initial value problem for compressible Euler systems and for Dafermos’s nonlinear oscillator [8]. For fixed final time $t_{1}$, we have:
Definition 2.1 (LAAP).
In a time interval $[t_{0},t_{1}]$, where $t_{1}\in{\mathbb{R}}$ is less than or equal to the infimum of the maximal time of existence for solutions in $S$, a solution ${\boldsymbol{u}}\in S$ is LAAP-admissible if the action $A({\boldsymbol{u}})(t_{0},t_{1})$ is not greater than the action of all other solutions in $S$. The solution ${\boldsymbol{u}}\in S$ is strictly LAAP-admissible when the inequality is strict.
In the context of the wild solutions of the Riemann problem of [2], in a number of cases LAAP identifies the 2-shock solution as the only strictly LAAP-admissible solution for all $t_{1}>t_{0}$. But if the set $S$ includes solutions other than the wild solutions constructed in [2] and the 2-shock solution, LAAP, as just explained, can be ineffective. Obviously, the action depends upon the value chosen for $t_{1}$. In fact, as discussed in the next Section, this is the main feature in the construction of Markfelder and Pellhammer [14]: some choices of $t_{1}$ allow the one-dimensional 2-shock solution to be LAAP-admissible, while others do not.
Here we formulate a local in time version of LAAP, called LAAP0, for which the 2-shock solution is admissible in the case of Riemann problems treated in [2], when the set $S$ also includes wild solutions and the hybrid solutions of [14]. To define LAAP0 we need the following definitions. (Note that in Definition 2.2 the time interval $t_{1}({\boldsymbol{v}})$ depends on the solution ${\boldsymbol{v}}$.)
Definition 2.2.
(i) Given a class $S$ of solutions, a solution ${\boldsymbol{u}}\in S$ is preferred to a solution ${\boldsymbol{v}}\in S$, ${\boldsymbol{v}}\neq{\boldsymbol{u}}$, under LAAP0 when there exists a time $t_{1}=t_{1}({\boldsymbol{v}})>t_{0}$ such that $A({\boldsymbol{u}})(t_{0},\,t)\leq A({\boldsymbol{v}})(t_{0},\,t)$ for all $t\in(t_{0},\,t_{1})$. When, in addition, $A({\boldsymbol{u}})(t_{0},\,t)<A({\boldsymbol{v}})(t_{0},\,t)$ for some $t\in(t_{0},\,t_{1})$, then ${\boldsymbol{u}}$ is strictly preferred to ${\boldsymbol{v}}$.
(ii) A solution ${\boldsymbol{u}}\in S$ is (strictly) LAAP0 admissible in $S$ when for every ${\boldsymbol{v}}\in S,\,{\boldsymbol{v}}\neq{\boldsymbol{u}},$ ${\boldsymbol{u}}$ is (strictly) preferred under LAAP0.
Remark 2.1.
In applications to weak solutions to systems of conservation laws the set $S$ is always to be included in the set of entropic weak solutions with the same initial data as in [9]. A unique entropic weak solution is trivially LAAP0 admissible as the comparison set $S$ becomes a singleton.
When $t\rightarrow A({\boldsymbol{u}})(t_{0},\,t)$ is differentiable to sufficient order from the right at $t=t_{0}$, we can let $t\rightarrow t_{0}^{+}$ and avoid dependence upon $t_{1}({\boldsymbol{v}})$ by defining two closely related rate criteria. We have:
Definition 2.3.
(i) ${\boldsymbol{u}}\in S$ satisfies the Action Rate Admissibility Criterion (ARAC) when for all ${\boldsymbol{v}}\in S$, ${\boldsymbol{v}}\neq{\boldsymbol{u}}$,
$\frac{d}{dt}A({\boldsymbol{u}})(t_{0},\,t)|_{t=t^{+}_{0}}\leq\frac{d}{dt}A({\boldsymbol{v}})(t_{0},\,t)|_{t=t^{+}_{0}}.$ (2.2)
(ii) ${\boldsymbol{u}}\in S$ satisfies the Strict Action Rate Admissibility criterion (sARAC) when for all ${\boldsymbol{v}}\in S$, ${\boldsymbol{v}}\neq{\boldsymbol{u}}$, there exists $k\in\mathbb{N}$ with $\frac{d^{r}}{dt^{r}}A({\boldsymbol{u}})(t_{0},\,t)|_{t=t^{+}_{0}}=\frac{d^{r}}{dt^{r}}A({\boldsymbol{v}})(t_{0},\,t)|_{t=t^{+}_{0}}$ for $r=0,\dots,k-1$ and
$\frac{d^{k}}{dt^{k}}A({\boldsymbol{u}})(t_{0},\,t)|_{t=t^{+}_{0}}<\frac{d^{k}}{dt^{k}}A({\boldsymbol{v}})(t_{0},\,t)|_{t=t^{+}_{0}}.$ (2.3)
Remark 2.2.
1. The action rate admissibility criteria of the previous definitions (when $k=1$) are identical in form to Dafermos’s entropy rate admissibility criterion [3] on replacing action by entropy. For barotropic fluids considered in this paper entropy is just the energy. When $k>1$ the strict action rate criterion can be analogously carried over to define a $k$-th order entropy rate criterion.
2. For general solutions, a definition of entropy rate is presented in [7, Sects 1.2,1.3] when the time derivative only exists almost everywhere. The corresponding general definition of the action rate is obtained by replacing the energy $E$ by the action $A$ in the discussion of [7].
Since by assumption $A({\boldsymbol{u}})(t_{0},\,t_{0})=A({\boldsymbol{v}})(t_{0},\,t_{0})=0$, the required relationships easily follow from Definition 2.2 and Definition 2.3. We obtain:
Proposition 2.1.
  1. 1.
    A LAAP0-admissible solution is ARAC-admissible.
  2. 2.
    An sARAC-admissible solution is strictly LAAP0-admissible.
We apply the principles defined in Propositions 2.2 and 2.3 to the two-dimensional Riemann problem for compressible isentropic Euler equation studied in e.g., [2, 9, 14].
3 The Riemann problem
This Section defines the Riemann problem to we wish to analyse and. for convenience, summarises properties of various solutions to be studied subsequently. First, however, elements of convex integration are briefly reviewed.
3.1 Convex integration
The two main elements of convex integration pertinent to the present study comprise the concepts of a sub-solution and corrugation. For equation (2.1) a sub-solution is the function ${\boldsymbol{w}}$ that in analogy with a sub-harmonic function satisfies the inequality
$\frac{d{\boldsymbol{w}}}{dt}-F({\boldsymbol{w}})\leq 0.$
Corrugation is a method of successively superposing on the sub-solution a sequence of oscillations at diminishing scale that in the limit converges to an appropriately defined (weak) exact solution to the equation of interest in the sense of distributions. In Nash’s historic paper [15] on isometric embedding of Riemannian manifolds, sub-solutions correspond to short embeddings. The challenge was to produce the corrugation algorithm which he termed stages. Gromov [11, 12] developed the technique while De Lellis and Székelyhidi [4, 5] applied convex integration to construct wild solutions to the Euler equations. The lecture notes by Markfelder [13] may be consulted for details. An overview is provided in the survey by De Lellis and Székelyhidi [6].
Chiodaroli and Kreml [2] apply the convex integration procedure to the 2-dimensional Riemann problem we are dealing with and, specifically, construct relevant sub-solutions. Hence we need to explain their construction to make it easier to understand our application of LAAP0 to the isentropic form of two dimensional barotropic Euler system discussed in [2].
3.2 Isentropic compressible Euler system
In what follows, the independent variables are time $t$ and position ${\boldsymbol{x}}$ such that $(t,{\boldsymbol{x}})\in[0,\infty)\times{\mathbb{R}}^{2}$. Let $\rho$ be the unknown mass density, ${\boldsymbol{m}}$ the momentum, and $p(\rho)$ the constitutively determined pressure. The fluid velocity is ${\boldsymbol{v}}={\boldsymbol{m}}/\rho.$ In an obvious notation, the two dimensional compressible barotropic Euler system of gas dynamics is given by
$\partial_{t}\rho+div_{x}{\boldsymbol{m}}$ $=$ $0,$ (3.1)
$\partial_{t}{\boldsymbol{m}}+div_{x}\left(\frac{{\boldsymbol{m}}\otimes{\boldsymbol{m}}}{\rho}\right)+\nabla_{x}p(\rho)$ $=$ $0,$ (3.2)
to which are adjoined initial conditions for $(\rho,{\boldsymbol{v}})$:
$(\rho,\,{\boldsymbol{v}})(0,\,.)=(\rho_{0},\,{\boldsymbol{v}}_{0}).$ (3.3)
The pair $(\rho,\,{\boldsymbol{v}})$ is a weak solution to the initial value problem for the barotropic compressible Euler system when the pair satisfy (3.1)-(3.3) in the sense of distributions.
The set $S$ of solutions to the initial value problem (3.1)-(3.3) consists of all entropic weak solutions, that is, weak solutions that satisfy the energy-entropy condition in the sense of distributions (see for example [2, 3, 13, 14]):
$\partial_{t}\left(\rho\varepsilon(\rho)+\rho\frac{|{\boldsymbol{v}}|^{2}}{2}\right)+div_{x}\left[\left(\rho\varepsilon(\rho)+\rho\frac{|{\boldsymbol{v}}|^{2}}{2}\right){\boldsymbol{v}}\right]\leq 0,$ (3.4)
where $\varepsilon$, the specific internal energy, satisfies $\varepsilon^{\prime}(\rho)=p(\rho)/\rho^{2}.$ Throughout this paper it is assumed that $p^{\prime}(\rho)>0.$
Chiodaroli and Kreml [2] study the isentropic form of the barotropic system (3.1) and (3.2) for which $p(\rho)=\rho^{\gamma}$ with constant $\gamma\geq 1$. The Riemann initial data is
$(\rho_{0},\,{\boldsymbol{m}}_{0})=\left\{\begin{array}[]{lll}(\rho_{-},\,{\boldsymbol{m}}_{-}),&x_{2}<0,&-\infty<x_{1}<\infty,\\ (\rho_{+},\,{\boldsymbol{m}}_{+}),&x_{2}>0,&-\infty<x_{1}<\infty,\end{array}\right.$ (3.5)
where $\rho_{\pm},\,{\boldsymbol{m}}_{\pm}$ are constants.
The article by Tartar [16] is crucial for the next lemma [2, Lemma 3.2] which is here stated without proof. The lemma not only presents a method for constructing corrugations but also establishes that the corrugations are non-unique.
We denote by $S_{0}^{2\times 2}$ the set of $2\times 2$ symmetric matrices with zero trace.
Lemma 3.1.
Let $(\tilde{{\boldsymbol{v}}},\,\tilde{{\boldsymbol{u}}})\in\mathrm{I\!R\!}^{2}\times S^{2\times 2}_{0}$ and let $C>0$ be a positive constant such that
$\tilde{{\boldsymbol{v}}}\otimes\tilde{{\boldsymbol{v}}}-\tilde{{\boldsymbol{u}}}<\frac{C}{2}Id.$ (3.6)
For any $\Lambda\subset\mathrm{I\!R\!}^{+}\times\mathrm{I\!R\!}^{2}$ there are infinitely many maps $({\boldsymbol{v}},\,{\boldsymbol{u}})$ with the properties:
${\boldsymbol{v}}$ and ${\boldsymbol{u}}$ vanish identically outside $\Lambda$.
$div_{x}{\boldsymbol{v}}=0,\quad\partial_{t}{\boldsymbol{v}}+div_{x}{\boldsymbol{u}}=0,$ in the sense of distributions.
$(\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}})\otimes(\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}})-(\tilde{{\boldsymbol{u}}}+{\boldsymbol{u}})=\frac{C}{2}Id$  a.e. in $\Lambda$.
Several observations may be listed.
Remark 3.1 (Remarks on Lemma 3.1).
  1. 1.
    Lemma 3.1(i) states that the maps $({\boldsymbol{v}},\,{\boldsymbol{u}})$ have compact support $\Lambda$. In particular, these maps are restricted to a compact time interval and vanish smoothly outside this interval. Markfelder and Pellhammer [14] exploited the compactness when constructing a counterexample to the claim in [9]. Additional comment is provided in Section 5.
  2. 2.
    On appeal to [2, Sect. 6], the solutions $({\boldsymbol{v}},\,{\boldsymbol{u}})$ to the “pressureless” gas dynamics system of Lemma 3.1$(ii)$ may be taken as periodic in $x_{1}$. For later reference, we let the period be $2L_{3}$.
  3. 3.
    The reason for non-uniqueness of the convex integration solutions originates in the assertion in Lemma 3.1 that there are an infinite number of functions $({\boldsymbol{v}},\,{\boldsymbol{u}})$
  4. 4.
    The incompressible pressureless gas dynamics system specified in Lemma 3.1$(ii)$ is time reversible as proved on replacing $(t,\,x_{1},\,x_{2})$ by $(-t,\,-x_{1},\,-x_{2})$. Also, shocks cannot form in contrast to the usual time irreversible gas dynamics with non-zero pressure.
  5. 5.
    Another important implication of the last Remark is that “entropy” admissibility criteria may not always serve as appropriate selection principles. This obvious but non-trivial observation is the motivation not only for the present investigation but also for our earlier paper [9].
The next task is to define the sub-solution of Chiodaroli and Kreml [2, Sect. 3]. Together with the corrugations of Lemma 3.1 it is an essential component of the convex integration process.
A fan partition of $\mathrm{I\!R\!}^{2}\times(0,\,\infty)$ consists of the open sets
$P_{-}$ $=$ $\left\{(x_{1},\,x_{2},\,t):t>0,\quad x_{2}<\nu_{-}t,\quad-\infty<x_{1}<\infty\right\},$ (3.7)
$P_{1}$ $=$ $\left\{(x_{1},\,x_{2},\,t):t>0,\,\nu_{-}t<x_{2}<\nu_{+}t,\,-\infty<x_{1}<\infty\right\},$ (3.8)
$P_{+}$ $=$ $\left\{(x_{1},\,x_{2},\,t):t>0,\,\nu_{+}t<x_{2},\,-\infty<x_{1}<\infty\right\},$ (3.9)
where $\nu_{-}<\nu_{+}$ are arbitrary real numbers that correspond to the shock speeds and which depend upon $\bar{\rho}$. They are written $(\nu_{-}(\bar{\rho}),\,\nu_{+}(\bar{\rho}))$ to emphasize this dependence.
Definition 3.1 (Fan sub-solution).
A fan sub-solution to (3.1) and (3.2) subject to initial data (3.5) is the triple $(\bar{\rho},\,\bar{{\boldsymbol{v}}},\,\bar{{\boldsymbol{u}}}):\mathrm{I\!R\!}^{2}\times(0,\,\infty)\rightarrow(\mathrm{I\!R\!}^{+},\,\mathrm{I\!R\!}^{2},\,S^{2\times 2}_{0})$ of piecewise constant functions subject to:
The region $\mathrm{I\!R\!}^{2}\times(0,\,\infty)$ can be decomposed into fan partitions $P_{-},\,P_{1},\,P_{+}$ such that
$(\bar{\rho},\,\bar{{\boldsymbol{v}}},\,\bar{{\boldsymbol{u}}})=(\rho_{-},\,{\boldsymbol{v}}_{-},\,{\boldsymbol{u}}_{-})\textup{I}_{P_{-}}+(\rho_{1},\,{\boldsymbol{v}}_{1},\,{\boldsymbol{u}}_{1})\textup{I}_{P_{1}}+(\rho_{+},\,{\boldsymbol{v}}_{+},\,{\boldsymbol{u}}_{+})\textup{I}_{P_{+}},$ (3.10)
where $\rho_{1},\,{\boldsymbol{v}}_{1},\,{\boldsymbol{u}}_{1}$ are constants with $\rho_{1}>0$ and where ${\boldsymbol{u}}_{-}={\boldsymbol{v}}_{-}\otimes{\boldsymbol{v}}_{-}-(1/2)|{\boldsymbol{v}}_{-}|^{2}{\boldsymbol{I}d}$, with a similar definition for ${\boldsymbol{u}}_{+}$.
There exists a positive constant $C$ such that
${\boldsymbol{v}}_{1}\otimes{\boldsymbol{v}}_{1}-{\boldsymbol{u}}_{1}<\frac{C}{2}{\boldsymbol{I}d}.$ (3.11)
The triple $(\bar{\rho},\,\bar{{\boldsymbol{v}}},\,\bar{{\boldsymbol{u}}})$ is a solution in the sense of distributions to the system
$\partial_{t}\bar{\rho}+div_{x}(\bar{\rho}\bar{{\boldsymbol{v}}})$ $=$ $0,$ (3.12)
$\partial_{t}(\bar{\rho}\bar{{\boldsymbol{v}}})+div_{x}(\bar{\rho}\bar{{\boldsymbol{u}}})+\nabla_{x}\left(p(\bar{\rho})+\frac{1}{2}\left(C\rho_{1}{\boldsymbol{I}d}_{P_{1}}+\bar{\rho}|\bar{{\boldsymbol{v}}}|^{2}{\boldsymbol{I}d}_{P_{+}\cup P_{-}}\right)\right)$ $=$ $0$ (3.13)
It may be easily concluded from condition $(iii)$ that in $P_{-}\cup P_{+}$ the triple $(\bar{\rho},\,\bar{{\boldsymbol{v}}},\,\bar{{\boldsymbol{u}}})$ is a weak solution to (3.1), (3.2) and (3.3). In $P_{1}$, however, the triple in the sense of distributions satisfies
$\partial_{t}(\bar{\rho}\bar{{\boldsymbol{v}}})+div_{x}(\bar{\rho}\bar{{\boldsymbol{u}}})+\nabla_{x}\left(p(\bar{\rho})+\frac{1}{2}C\rho_{1}\right)=0.$ (3.14)
To obtain convex integration solutions to (3.1), (3.2) and (3.5), Chiodaroli and Kreml [2] add the fan sub-solution to the maps $({\boldsymbol{v}},\,{\boldsymbol{u}})$ of Lemma 3.1. Hence, set $\Lambda=P_{1}$ and take $\tilde{{\boldsymbol{v}}}={\boldsymbol{v}}_{1},\,\tilde{{\boldsymbol{u}}}={\boldsymbol{u}}_{1}$ to obtain
$\partial_{t}(\rho_{1}({\boldsymbol{v}}+{\boldsymbol{v}}_{1}))=0.$ (3.15)
Furthermore, from (3.14) upon elimination of the constant $C$ using Lemma 3.1$(iii)$ we have in $P_{1}$:
$\partial_{t}(\rho_{1}({\boldsymbol{v}}+{\boldsymbol{v}}_{1}))+div_{x}(\rho_{1}({\boldsymbol{u}}+{\boldsymbol{u}}_{1}))+\nabla_{x}\left[p(\rho_{1})+\rho_{1}(({\boldsymbol{v}}+{\boldsymbol{v}}_{1})\otimes({\boldsymbol{v}}+{\boldsymbol{v}}_{1})-({\boldsymbol{u}}+{\boldsymbol{u}}_{1})\right]=0,$ (3.16)
where the trivial term from (3.15) has been retained. The second and last terms in (3.16) cancel and we conclude that $(\rho_{1},\,{\boldsymbol{v}}+{\boldsymbol{v}}_{1})$ satisfy (3.1) and (3.2) in $P_{1}$. An infinite number of weak solutions therefore has been obtained to the Riemann initial value problem with initial conditions (3.5). The corresponding shock speeds depend upon $\bar{\rho}=\rho_{1}$ and are expressed as $\nu_{-}(\rho_{1}),\,\nu_{+}(\rho_{1})$. Explicit expressions are presented in [2, (4.46),(4.47),(4.51),(4.52)]. Moreover, the trace of the expression given in Lemma 3.1 $(iii)$ when $\Lambda=P_{1}$ yields
$C=|\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}}|^{2}\textup{I}_{P_{1}}.$ (3.17)
On the other hand, Chiodaroli and Kreml [2] introduce functions of $\rho_{1}$, $\beta$, $\varepsilon_{1}$, and a parameter $\varepsilon_{2}$ such that $C$ defined in (3.17) can be written as
$C=\beta^{2}(\rho_{1})+\varepsilon_{1}(\rho_{1})+\varepsilon_{2}.$ (3.18)
Let $I$ denote some left interval of $\rho_{m}$ given by
$I=[\rho^{\ast},\,\rho_{m}].$ (3.19)
Then for $\rho_{1}\in I$, the functions $\beta(\rho_{1})$ and $\varepsilon_{1}(\rho_{1})$ are smooth, while $\varepsilon_{2}>0$. In particular,
$\beta^{2}(\rho_{m})=v^{2}_{m},\qquad\varepsilon_{1}(\rho_{m})=0,$ (3.20)
where $\rho_{m}$ and $v_{m}$ refer respectively to the density and second component of the velocity for the one-dimensional classical 2-shock solution to the Riemann initial value isentropic problem. We have for the 2-shock solution
$\left\{\begin{array}[]{llllll}\rho=\rho_{-},&t>0,&v_{a}=0,&v_{b}=v_{b_{-}},&-\infty<x_{1}<\infty,&x_{2}<\nu_{-}(\rho_{m})t,\\ \rho=\rho_{+},&t>0,&v_{a}=0,&v_{b}=v_{b_{+}},&-\infty<x_{1}<\infty,&x_{2}>\nu_{+}(\rho_{m})t,\\ \rho=\rho_{m},&t>0,&v_{a}=0,&v_{b}=v_{m},&-\infty<x_{1}<\infty,&\nu_{-}(\rho_{m})t<x_{2}<\nu_{+}(\rho_{m})t.\end{array}\right.$
Here, we have set ${\boldsymbol{v}}=(v_{a},\,v_{b})$ and note that $\nu_{-}(\rho_{m}),\,\nu_{+}(\rho_{m}),\,\rho_{-},\,\rho_{+},\,\rho_{m},\,v_{b_{-}},\,v_{b_{+}},,v_{m}$ are determined consistent with the Rankine-Hugoniot conditions.
The following sections apply the admissibility criterion LAAP0, formulated in Section 2, to the two-dimensional compressible isentropic Euler equations in the class $S$ consisting of the classical 2-shock solution, the convex integration solutions of [2] and the hybrid solutions of [14].
4 Justification for LAAP0
In summary, Chiodaroli and Kreml [2] treat the two-dimensional isentropic compressible Euler equation and for specific Riemann initial data prove that there are an infinite number of entropic weak convex integration solutions. For some sub-solutions, the wild solutions constructed in [2] dissipate energy at $t=0^{+}$ more rapidly than the one-dimensional 2-shock solution. As a result. for some sub-solution data the 2-shock solution is not selected by the entropy rate criterion.
In the same Riemann initial value problem, LAAP prefers the 2-shock solution to the entropic wild solutions constructed in [2]. However, Markfelder and Pellhammer [14] exploited Lemma 3.1 to construct examples of solutions which are preferred by LAAP to the 2-shock solution. To this end, note that Lemma 3.1$(i)$ restricts to a compact time interval the two dimensional component of the fan sub-solution, which therefore vanishes smoothly outside this interval. Consequently, for time $T_{0}$ greater than the upper limit of this compact time interval, a new problem may be constructed with piece-wise constant initial data determined by the fan sub-solution. The classical method of elementary waves is then used to extend locally in time the solution obtained in [2]. Hence Markfelder and Pellhammer [14] obtain hybrid solutions whose action for which for sufficiently large finite times is smaller than that of the 2-shock solution. It is worth remarking that the hybrid solution constructed by Markfelder and Pellhammer cannot be trivially extended to a global in time solution as the initial data at $t=T_{0}$ do not necessarily satisfy the small bounded variation hypothesis of Glimm’s global existence theorem [10].
These developments motivate a new admissibility criterion to replace the one formulated in [9]. The new feature is that the actions are defined with variable terminal time $t_{1}$ that enables the limit $t_{1}\rightarrow t_{0}^{+}$ to be taken.
5 Application of LAAP0
Recall that the Lagrangian $\mathcal{L}$ of fluid motion is the pointwise difference between the kinetic and potential energy and for the barotropic compressible fluid is given by
$\mathcal{L}=\frac{1}{2}\rho\left|{\boldsymbol{v}}\right|^{2}-\rho\varepsilon(\rho),$ (5.1)
in the notation of (3.4).
The corresponding time-dependent action $A(\rho,\,{\boldsymbol{v}})$ in a domain $\Lambda=\Omega\times[t_{o},\,t]$ is given by
[…]
Put
[…]
and note that the 2-shock solution becomes identical to the convex integration solution of [2] in the region external to $\Lambda=P_{1}=[-L_{3},\,L_{3}]\times[\ell_{1},\,\ell_{2}]\times[0,\,t]$ where $2L_{3}$ is defined in Remark 3.1$(ii)$, the left interval $I$ is defined in (3.19),and
$\ell_{1}<\min{(\nu_{-}(\rho_{1}:\rho_{1}\in I)}t,\qquad\ell_{2}>\max{(\nu_{+}(\rho_{1}:\rho_{1}\in I)}t.$ (5.3)
See Figure 1. In consequence, we have
Figure 1: Domains of integration
$L_{diff}(\rho_{1})=\frac{1}{2}\rho_{m}v^{2}_{m}-\rho_{m}\varepsilon(\rho_{m})-\frac{1}{2}\rho_{1}C+\rho_{1}\varepsilon(\rho_{1})$ (5.4)
where $C$ is given by (3.18). Set
[…]
The function $D(\rho_{1},\,t)$ is continuous with respect to $\rho_{1}\in I$. Choose $\rho_{1}=\rho_{m}$ to obtain
[…]
On noting from (5.4) that
$L_{diff}(\rho_{m})=-\frac{1}{2}\rho_{m}\varepsilon_{2}<0,$
we conclude that $D(\rho_{1},t)<0$. Further, because at $t=0$ the domain of integration shown in Figure 1, where the integrand is non-zero, collapses to a set of measure zero, we have
$\frac{\partial}{\partial t}D(\rho_{1},t)|_{t=0^{+}}=0.$ (5.5)
Finally, we note that
$\frac{\partial^{2}}{\partial t^{2}}D(\rho_{1},t)|_{t=0^{+}}<0$ (5.6)
in a sufficiently small left neighbourhood $I$ of $\rho_{m}$, because this inequality holds at $\rho_{1}=\rho_{m}$ and the second time derivative of $D(\rho_{1},t)$ is continuous in $\rho_{1}$.
We now state and prove the main results.
Consider the Riemann initial value problem for the two-dimensional isentropic compressible Euler system (3.1), (3.2) and (3.5). Let the set $S$ of solutions consist of the one-dimensional 2-shock solution, the entropic global convex integration weak solutions constructed by Chiodaroli and Kreml [2], and the hybrid solutions constructed by Markfelder and Pellhammer [14] with fan sub-solution data $\rho_{1}\in I$ and $\varepsilon_{2}>0$. Recall that for small $t>0$ the hybrid solutions coincide with those of [2]. We have
Theorem 5.1.
Let $\varepsilon_{2}>0$ and let $\rho_{1}$ belong to a fixed left interval $I$ of $\rho_{m}$ so that a fan sub-solution is defined. Then the 2-shock solution is the sARAC-, and therefore the strictly LAAP0-admissible solution in $S$.
Proof.
The proof employs components of the proof of [9, Thm. 5], combined with Definitions 2.2 and 2.3. ∎
As just shown, for $\rho_{1}$ in the interval $I$ the action and its first and second time derivatives are lower for the 2-shock solution than for any other solution in $S$, including those obtained in [14] since for small enough time these coincide with those of [2]. Hence, both sARAC and strict LAAP0 exclude the admissibility of the convex integration and hybrid solutions and establish the strict admissibility of the 2-shock solution in $S$.
[…]
Theorem 5.2.
Let $\gamma=2.$ For any Riemann data satisfying the necessary condition for the existence of the classical 2-shock solution given by
$(v_{b-}-v_{b+})^{2}\rho_{+}\rho_{-}-(\rho_{+}-\rho_{-})(p(\rho_{+}-p(\rho_{-}))>0,$ (5.7)
and any $\varepsilon_{2}>0$ in (3.18), the classical 2-shock solution is $sARAC$- and strictly LAAP0-admissible in $S$.
Proof.
The proof is a straightforward adaptation of [9, Thm. 6] and Theorem 5.1. ∎
Acknowledgements
The authors are grateful for support from the ICMS Research in Groups scheme. They also thank C. M. Dafermos, B. Gebhard and O. Kreml for valuable remarks.