### v5

# On action rate admissibility criteria

Abstract
We formulate new admissibility criteria for initial value problems motivated by the least action principle. These are applied to a two-dimensional Riemann initial value problem for the isentropic compressible Euler fluid flow. It is shown that the criterion prefers the 2-shock solution to solutions obtained by convex integration by Chiodaroli and Kreml or to the hybrid solutions recently constructed by Markfelder and Pellhammer.
MSC: 35L65 (primary), 35C06, 35D30, 35LQ35, 76N10 (secondary).
Keywords: Weak Solution; Riemann Problem; Least Action Principle; Convex Integration ; Admissible Solution; Entropy Rate.
1 Introduction
The paper [9] introduces a version of the least action principle as a selection criterion in initial value problems possessing non-unique solutions. Interest arises from examples that use convex integration techniques to obtain families of non-unique wild solutions to the Euler system (De Lellis and Székelyhidi [4, 5]), and non-unique solutions to the Navier-Stokes equations (Buckmaster and Vicol [1]). Convex integration is also employed by Chiodaroli and Kreml [2] to investigate the two-dimensional Riemann initial value problem for the isentropic compressible Euler system. They construct a large set of entropic wild solutions to the Riemann problem. Consequently, in these problems it is desirable to look for an admissibility criterion that distinguishes a particular solution from the many possible. Because of its mathematical simplicity, we are motivated to seek an admissibility criterion that selects the classical 2-shock solution.
Other admissibility criteria in fluid dynamics include the principles of minimum potential energy or entropy, various rate of change criteria, and that of least action (LAAP) developed in [9]. The latter principle applied to the initial value problem considered in [2] selects the 2-shock solution provided the solutions obtained by convex integration are the only other ones admitted for comparison. On the other hand, Chiodaroli and Kreml [2] show that there exist Riemann data for which there are wild solutions that possess an entropy rate lower than the 2-shock solution which therefore is not selected by Dafermos’s entropy rate criterion [3]. Nevertheless, LAAP selects the 2-shock in preference to wild solutions in particular cases.
LAAP, however, does not select the 2-shock solution when the set of comparison solutions is enlarged to include other weak solutions in addition to the 2-shock and convex integration solutions. Markfelder and Pellhammer [14] recently constructed hybrid solutions to the same Riemann initial value problem for which, subject to prescribed initial data, LAAP does not identify the 2-shock solutions as being uniquely admissible. In fact, this result indicates that LAAP does not provide an admissible solution to the Riemann problem. Accordingly, an alternative to LAAP must be sought when the 2-shock solution in this Riemann problem is preferred.
The main purpose of this paper therefore is to formulate a new admissibility criterion, called LAAP0, which for the Riemann problem considered in [2] selects the 2-shock solution in comparison to the convex integration and hybrid solutions. LAAP0, while still motivated by the least action principle, involves a rate of change of action. Both the Dafermos entropy rate criterion and LAAP0 are local in time. The solutions constructed in [14] are excluded as they rely upon the global nature of LAAP.
Section 2 collects basic definitions, states the least action admissibility principle (LAAP) derived in [9] and formulates the new least action admissibility principle (LAAP0) and two closely related rate criteria. Section 3, after a brief description of convex integration, introduces the two-dimensional Riemann initial value problem for the barotropic and isentropic compressible Euler equations of gas dynamics and summarises further material from [2] including the construction of sub-solutions. The 2-shock solution is also described. Justification for the introduction of LAAP0 is provided in Section 4, while Section 5 defines the relevant notion of action and applies LAAP0 to the Riemann problem treated in [2] to establish the admissibility of the 2-shock solution.
Normally accepted notation is introduced without comment.
2 Least action admissibility criteria
Let $S$ be a set of solutions $\boldsymbol{u}$} of the Cauchy problem for the evolutionary system specified by
$\frac{d{\boldsymbol{u}}}{dt}=F({\boldsymbol{u}}),\qquad{\boldsymbol{u}}(t_{0})={\boldsymbol{u}}_{0},\qquad t\in(t_{0},\,T],$ (2.1)
where $F({\boldsymbol{u}})$ is a given operator and $t_{0},\,T]$ is the maximal time interval of existence of solutions belonging to $S$. Global existence is not assumed so that possibly $T<\infty$. Later, the system (2.1) is specialised to be the Euler-Lagrange equation derived from a least action principle.
Suppose that $S$ is not a singleton, which implies that (2.1) has non-unique solutions. The objective is to formulate admissibility criteria stated in terms of the action whose strict forms select unique elements of $S$. The definition of the action to be used is postponed to Section 5. For present purposes, it is sufficient to assume that for each solution ${\boldsymbol{u}}\in S$, the action $A({\boldsymbol{u}})(t_{0},\,t)$ vanishes at $t=t_{0}$ for all ${\boldsymbol{u}}\in S$.
To start, it is convenient to recall the Least Action Admissibility Principle (LAAP) introduced in [9] and applied to the Riemann initial value problem for compressible Euler systems and to Dafermos’ nonlinear oscillator [8]. For fixed final time $t_{1}$, we have:
Definition 2.1 (LAAP).
In a time interval $t_{0},t_{1}]$, where $t_{1}\in{\mathbb{R}}$ is less than or equal to the infimum of the maximal time of existence for solutions in $S$, a solution ${\boldsymbol{u}}\in S$ is LAAP-admissible if the action $A({\boldsymbol{u}})(t_{0},t_{1})$ is not greater than the action of all other solutions in $S$. The solution ${\boldsymbol{u}}\in S$ is strictly LAAP-admissible when the inequality is strict.
In the context of the wild solutions of the Riemann problem of [2], in a number of cases LAAP identifies the 2-shock solution as the only strictly LAAP-admissible solution for all $t_{1}>t_{0}$. But when the set $S$ includes solutions other than the wild solutions constructed in [2] and the 2-shock solution, LAAP, as explained in Section 1, can be ineffective. Obviously, the action depends upon the value chosen for $t_{1}$. In fact, as discussed in the next Section, this is the main feature in the construction of Markfelder and Pellhammer [14]: some choices of $t_{1}$ allow the one-dimensional 2-shock solution to be LAAP-admissible, while others do not.
Here we formulate a local in time version of LAAP, called LAAP0, for which the 2-shock solution is admissible in particular cases of the Riemann problem treated in [2]. It is equally admissible when the set $S$ is enlarged to include the hybrid solutions of [14].
To define LAAP0 we need the following definitions. (Note that in Definition 2.2 and subsequently the time $t_{1}({\boldsymbol{v}})$ depends upon the solution $\boldsymbol{v}$}, and $t_{0}$ is such that $t_{0},T]$ is the maximal existence interval.)
Definition 2.2.
(i) Given a class $S$ of solutions, a solution ${\boldsymbol{u}}\in S$ is preferred to a solution ${\boldsymbol{v}}\in S$, ${\boldsymbol{v}}\neq{\boldsymbol{u}}$, under LAAP0 when there exists a time $t_{1}=t_{1}({\boldsymbol{v}})>t_{0}$ such that $A({\boldsymbol{u}})(t_{0},\,t)\leq A({\boldsymbol{v}})(t_{0},\,t)$ for all $t\in(t_{0},\,t_{1})$. When, in addition, $A({\boldsymbol{u}})(t_{0},\,t)<A({\boldsymbol{v}})(t_{0},\,t)$ for some $t\in(t_{0},\,t_{1})$, then $\boldsymbol{u}$} is strictly preferred to $\boldsymbol{v}$}.
(ii) A solution ${\boldsymbol{u}}\in S$ is (strictly) LAAP0 admissible in $S$ when for every ${\boldsymbol{v}}\in S,\,{\boldsymbol{v}}\neq{\boldsymbol{u}},$ $\boldsymbol{u}$} is (strictly) preferred under LAAP0.
Remark 2.1.
In applications to weak solutions to systems of conservation laws the set $S$ is always to be included in the set of entropic weak solutions with the same initial data as in [9]. A unique entropic weak solution is trivially LAAP0 admissible as the comparison set $S$ becomes a singleton.
When $t\rightarrow A({\boldsymbol{u}})(t_{0},\,t)$ is differentiable to sufficient order from the right at $t=t_{0}$, we can let $t\rightarrow t_{0}^{+}$ and avoid dependence upon $t_{1}({\boldsymbol{v}})$ by defining two closely related rate criteria. We have:
Definition 2.3.
(i) ${\boldsymbol{u}}\in S$ satisfies the Action Rate Admissibility Criterion (ARAC) when for all ${\boldsymbol{v}}\in S$, ${\boldsymbol{v}}\neq{\boldsymbol{u}}$,
$\frac{d}{dt}A({\boldsymbol{u}})(t_{0},\,t)|_{t=t^{+}_{0}}\leq\frac{d}{dt}A({\boldsymbol{v}})(t_{0},\,t)|_{t=t^{+}_{0}}$. (2.2)
(ii) ${\boldsymbol{u}}\in S$ satisfies the Strict Action Rate Admissibility criterion (sARAC) when for all ${\boldsymbol{v}}\in S$, ${\boldsymbol{v}}\neq{\boldsymbol{u}}$, there exists $k\in\mathbb{N}$ with $\frac{d^{r}}{dt^{r}}A({\boldsymbol{u}})(t_{0},\,t)|_{t=t^{+}_{0}}=\frac{d^{r}}{dt^{r}}A({\boldsymbol{v}})(t_{0},\,t)|_{t=t^{+}_{0}}$ for $r=0,\dots,k-1$ and
$\frac{d^{k}}{dt^{k}}A({\boldsymbol{u}})(t_{0},\,t)|_{t=t^{+}_{0}}<\frac{d^{k}}{dt^{k}}A({\boldsymbol{v}})(t_{0},\,t)|_{t=t^{+}_{0}}$. (2.3)
Remark 2.2.
1. The action rate admissibility criteria of the previous definitions (when $k=1$) are identical in form to Dafermos’s entropy rate admissibility criterion [3] on replacing action by entropy. For barotropic fluids considered in this paper entropy is just the energy. When $k>1$ the strict action rate criterion can be analogously carried over to define a $k$-th order entropy rate criterion.
2. For a class of solutions more general than considered here, a definition of entropy rate is presented in [7, Sects 1.2,1.3] when the time derivative only exists almost everywhere. The corresponding general definition of the action rate is obtained by replacing the energy $E$ by the action $A$ in the discussion of [7].
Since by assumption $A({\boldsymbol{u}})(t_{0},\,t_{0})=A({\boldsymbol{v}})(t_{0},\,t_{0})=0$, the required relationships easily follow from Definition 2.2 and Definition 2.3. We obtain:
Proposition 2.1.
  1. 1.
    A LAAP0-admissible solution is ARAC-admissible.
  2. 2.
    An sARAC-admissible solution is strictly LAAP0-admissible.
We next apply the principles introduced in Definitions 2.2 and 2.3 to the two-dimensional Riemann problem for compressible isentropic Euler equation studied in e.g., [2, 9, 14].
3 The Riemann problem for a compressible barotropic Euler system
This Section defines the 2-dimensional Riemann problem and the isentropic form to be analysed. For convenience, several subsections are devoted to summarising various solutions and their relevant properties required in the later application of LAAP0. In particular, we consider solutions constructed by Chiodaroli and Kreml [2] using convex integration techniques, the two main elements of which are first briefly reviewed.
3.1 Convex integration
The two main elements of the convex integration procedure pertinent to the present study are the concepts of a sub-solution and of corrugation. For equation (2.1) a sub-solution is the function $\boldsymbol{w}$} that satisfies the inequality
$\frac{d{\boldsymbol{w}}}{dt}-F({\boldsymbol{w}})\leq 0.$
By corrugation we mean an iterative scheme of superposing on the sub-solution a sequence of spatio-temporal oscillations at decreasingly small scales such that in the limit the combined subsolution and oscillation converge to an appropriately defined weak (exact) solution to the equation of interest. In Nash’s fundamental paper [15] on isometric embedding of Riemannian manifolds, sub-solutions correspond to short embeddings. Nash produced a corrugation algorithm which he termed stages. Gromov [11, 12] generalised the technique, while De Lellis and Székelyhidi [4, 5] applied the general theory of convex integration to construct wild solutions to the Euler equations. The lecture notes by Markfelder [13] may be consulted for details. An overview is provided in the survey by De Lellis and Székelyhidi [6].
Chiodaroli and Kreml [2] apply the convex integration procedure to the 2-dimensional Riemann problem for the isentropic compressible Euler system in the case when it admits a 2-shock solution. We briefly explain their construction of sub-solutions and of weak solutions and refer to the latter as wild solutions.
3.2 Isentropic compressible Euler system
In what follows, the independent variables are time $t$ and position $\boldsymbol{x}$} such that $(t,{\boldsymbol{x}})\in[0,\infty)\times{\mathbb{R}}^{2}$. Let $\rho>0$ be the unknown mass density, $\boldsymbol{v}$} the fluid velocity, and $p(\rho)$ the constitutively defined pressure. In an obvious notation, the two dimensional compressible barotropic Euler system of gas dynamics is given by
$\partial_{t}\rho+div_{x}(\rho{\boldsymbol{v}})$ $=$ $0,$ (3.1)
$\partial_{t}(\rho{\boldsymbol{v}})+div_{x}\left(\rho{\boldsymbol{v}}\otimes{\boldsymbol{v}}\right)+\nabla_{x}[p(\rho)]$ $=$ $0,$ (3.2)
to which are adjoined the initial conditions
$(\rho,\,{\boldsymbol{v}})(0,\,\cdot)=(\rho_{0},\,{\boldsymbol{v}}_{0}).$ (3.3)
The pair $(\rho,\,{\boldsymbol{v}})$ is a weak solution to the initial value problem for the system (3.1)-(3.3) when it satisfies this system in the sense of distributions.
We consider entropic weak solutions to the initial value problem (3.1)-(3.3), that is, weak solutions that satisfy the energy-entropy condition in the sense of distributions (see for example [2, 3, 13, 14]):
$\partial_{t}\left(\rho\varepsilon(\rho)+\rho\frac{|{\boldsymbol{v}}|^{2}}{2}\right)+div_{x}\left[\left(\rho\varepsilon(\rho)+\rho\frac{|{\boldsymbol{v}}|^{2}}{2}\right){\boldsymbol{v}}\right]\leq 0,$ (3.4)
where $\varepsilon$, the specific internal energy, satisfies $\varepsilon^{\prime}(\rho)=p(\rho)/\rho^{2}.$ Throughout this paper it is assumed that $p^{\prime}(\rho)>0.$
Chiodaroli and Kreml [2] study the isentropic form of the barotropic system (3.1) and (3.2) for which $p(\rho)=\rho^{\gamma}$, where the constant $\gamma$ satisfies $1<\gamma\leq 3$. The Riemann initial data is
$(\rho_{0},\,{{\boldsymbol{v}}_{0}})=\left\{\begin{array}[]{lll}(\rho_{-},\,{{\boldsymbol{v}}_{-}}),&x_{2}<0,&-\infty<x_{1}<\infty,\\ (\rho_{+},\,{{\boldsymbol{v}}_{+}}),&x_{2}>0,&-\infty<x_{1}<\infty,\end{array}\right$. (3.5)
where $\rho_{\pm},\,{{\boldsymbol{v}}_{\pm}}$ are constants.
In [2, (2.47), Lemma 2.4(5)], Chiodaroli and Kreml show that for any Riemann data satisfying the necessary condition
$(v_{-}-v_{+})^{2}\rho_{+}\rho_{-}-(\rho_{+}-\rho_{-})(p(\rho_{+})-p(\rho_{-}))>0,$ (3.6)
there exists a unique self-similar solution to the initial value problem, the 2-shock solution, given by
$\left\{\begin{array}[]{lllll}\rho=\rho_{-},&v_{b}=v_{{-}},&t>0,&-\infty<x_{1}<\infty,&x_{2}<\nu_{-}t,\\ \rho=\rho_{+},&v_{b}=v_{{+}},&t>0,&-\infty<x_{1}<\infty,&x_{2}>\nu_{+}t,\\ \rho=\rho_{m},&v_{b}=v_{m},&t>0,&-\infty<x_{1}<\infty,&\nu_{-}t<x_{2}<\nu_{+}t.\end{array}\right$.
Here ${\boldsymbol{v}}=(0,\,v_{b})$; $\rho_{m}$ and $v_{m}$ are found from [2, (2.48), (2.49)] and the speeds of the shocks $\nu_{\pm}$ are determined from the Riemann data by the Rankine-Hugoniot conditions; see [2, (2.29), (2.32)].
3.3 Fan sub-solutions
The next task is to define the sub-solution of Chiodaroli and Kreml [2, Sect. 3]. Together with the corrugations of Lemma 3.1 it is an essential component of obtaining weak solutions by the convex integration process.
A fan partition of $\mathrm{I\!R\!}^{2}\times(0,\,\infty)$ consists of the open sets
$P_{-}$ $=$ $\left\{(x_{1},\,x_{2},\,t):t>0,\quad x_{2}<\nu_{-}t,\quad-\infty<x_{1}<\infty\right\},$ (3.7)
$P_{1}$ $=$ $\left\{(x_{1},\,x_{2},\,t):t>0,\,\nu_{-}t<x_{2}<\nu_{+}t,\,-\infty<x_{1}<\infty\right\},$ (3.8)
$P_{+}$ $=$ $\left\{(x_{1},\,x_{2},\,t):t>0,\,\nu_{+}t<x_{2},\,-\infty<x_{1}<\infty\right\},$ (3.9)
where $\nu_{-}<\nu_{+}$ are arbitrary real numbers that correspond to the shock speeds.
We denote by $S_{0}^{2\times 2}$ the set of $2\times 2$ symmetric matrices with zero trace.
Definition 3.1 (Fan sub-solution).
A fan sub-solution to (3.1) and (3.2) subject to initial data (3.5) is the triple $(\bar{\rho},\,\bar{{\boldsymbol{v}}},\,\bar{{\boldsymbol{u}}}):\mathrm{I\!R\!}^{2}\times(0,\,\infty)\rightarrow(\mathrm{I\!R\!}^{+},\,\mathrm{I\!R\!}^{2},\,S^{2\times 2}_{0})$ of piecewise constant functions subject to:
The region $\mathrm{I\!R\!}^{2}\times(0,\,\infty)$ can be decomposed into fan partitions $P_{-},\,P_{1},\,P_{+}$ such that
$(\bar{\rho},\,\bar{{\boldsymbol{v}}},\,\bar{{\boldsymbol{u}}})=(\rho_{-},\,{\boldsymbol{v}}_{-},\,{\boldsymbol{u}}_{-})\textup{I}_{P_{-}}+(\rho_{1},\,{\boldsymbol{v}}_{1},\,{\boldsymbol{u}}_{1})\textup{I}_{P_{1}}+(\rho_{+},\,{\boldsymbol{v}}_{+},\,{\boldsymbol{u}}_{+})\textup{I}_{P_{+}},$ (3.10)
where $\rho_{1},\,{\boldsymbol{v}}_{1},\,{\boldsymbol{u}}_{1}$ are constants with $\rho_{1}>0$ and where ${\boldsymbol{u}}_{-}={\boldsymbol{v}}_{-}\otimes{\boldsymbol{v}}_{-}-(1/2)|{\boldsymbol{v}}_{-}|^{2}\mathrm{Id}$, with a similar definition for $\boldsymbol{u}$$_{+}$.
There exists a positive constant $C$ such that
${\boldsymbol{v}}_{1}\otimes{\boldsymbol{v}}_{1}-{\boldsymbol{u}}_{1}<\frac{C}{2}\mathrm{Id}.$ (3.11)
The triple $(\bar{\rho},\,\bar{{\boldsymbol{v}}},\,\bar{{\boldsymbol{u}}})$ is a solution in the sense of distributions to the system
$\partial_{t}\bar{\rho}+div_{x}(\bar{\rho}\bar{{\boldsymbol{v}}})$ $=$ $0,$ (3.12)
$\partial_{t}(\bar{\rho}\bar{{\boldsymbol{v}}})+div_{x}(\bar{\rho}\bar{{\boldsymbol{u}}})+\nabla_{x}\left(p(\bar{\rho})+\frac{1}{2}\left(C\rho_{1}\textup{I}_{P_{1}}+\bar{\rho}|\bar{{\boldsymbol{v}}}|^{2}\textup{I}_{P_{+}\cup P_{-}}\right)\right)$ $=$ $0.$ (3.13)
It may be easily concluded from condition $(iii)$ that in $P_{-}\cup P_{+}$ the triple $(\bar{\rho},\,\bar{{\boldsymbol{v}}},\,\bar{{\boldsymbol{u}}})$ is a weak solution to (3.1), (3.2) and (3.3). In $P_{1}$, however, the triple in the sense of distributions satisfies
$\partial_{t}(\bar{\rho}\bar{{\boldsymbol{v}}})+div_{x}(\bar{\rho}\bar{{\boldsymbol{u}}})+\nabla_{x}\left(p(\bar{\rho})+\frac{1}{2}C\rho_{1}\right)=0.$ (3.14)
3.4 Corrugations
In their article Chiodaroli and Kreml [2] prove non-uniqueness of weak solutions to the Riemann initial value problem with initial conditions (3.5), i.e. not only is there the classical 2-shock solution given in Section 3.2, but also an infinite number of weak solutions that possess non-trivial $x_{1}$ dependence. Their proof, based upon convex integration, consists of two pieces: (i) a base sub-solution and (ii) an infinite sequence of corrugations which when combined with (i) yields an infinite number of solutions to the initial value problem. Here we recall these two concepts within the context of our Riemann initial value problem for the isentropic Euler equations. We emphasise that the corrugations combined with even a single sub-solution yield an infinite number of weak solutions to initial value problem with initial conditions (3.5).
The article by Tartar [16] is crucial for the next lemma [2, Lemma 3.2] which is here stated without proof. The lemma not only presents a method for constructing corrugations but also establishes that the corrugations are non-unique.
Lemma 3.1.
Let $(\tilde{{\boldsymbol{v}}},\,\tilde{{\boldsymbol{u}}})\in\mathrm{I\!R\!}^{2}\times S^{2\times 2}_{0}$ and let $C>0$ be a positive constant such that
$\tilde{{\boldsymbol{v}}}\otimes\tilde{{\boldsymbol{v}}}-\tilde{{\boldsymbol{u}}}<\frac{C}{2}\mathrm{Id}.$ (3.15)
For any $\Lambda\subset\mathrm{I\!R\!}^{+}\times\mathrm{I\!R\!}^{2}$ there are infinitely many maps $({\boldsymbol{v}},\,{\boldsymbol{u}})$ with the properties:
$\boldsymbol{v}$} and $\boldsymbol{u}$} vanish identically outside $\Lambda$.
$div_{x}{\boldsymbol{v}}=0,\quad\partial_{t}{\boldsymbol{v}}+div_{x}{\boldsymbol{u}}=0,$ in the sense of distributions.
$(\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}})\otimes(\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}})-(\tilde{{\boldsymbol{u}}}+{\boldsymbol{u}})=\frac{C}{2}\mathrm{Id}$  a.e. in $\Lambda$.
Several observations may be listed.
Remark 3.1 (Remarks on Lemma 3.1).
  1. 1.
    Markfelder and Pellhammer [14] exploited in their discussion that $({\boldsymbol{v}},\,{\boldsymbol{u}})$ of Lemma 3.1 vanish after a finite time, allowing them to construct a hybrid weak solution. Additional comment is provided in Section 5.
  2. 2.
    On appeal to [2, Sect. 6], the solutions $({\boldsymbol{v}},\,{\boldsymbol{u}})$ to the system of Lemma 3.1 may be taken as periodic in $x_{1}$. For later reference, we let the period be $2L_{3}$.
  3. 3.
    The reason for non-uniqueness of the convex integration solutions originates in the assertion in Lemma 3.1 that there are an infinite number of functions $({\boldsymbol{v}},\,{\boldsymbol{u}})$ .
  4. 4.
    The system specified in Lemma 3.1 is time reversible as proved on replacing $(t,\,x_{1},\,x_{2})$ by $(-t,\,-x_{1},\,-x_{2})$.
  5. 5.
    Another important implication of the last Remark is that “entropy” admissibility criteria may not always serve as appropriate selection principles. This obvious but non-trivial observation is the motivation not only for the present investigation but also for our earlier paper [9].
3.5 Convex integration solutions
To obtain convex integration solutions to (3.1), (3.2) and (3.5), Chiodaroli and Kreml [2, Proposition 3.1] add the maps $({\boldsymbol{v}},\,{\boldsymbol{u}})$ of Lemma 3.1 to a fan sub-solution. Hence, set $\Lambda=P_{1}$ and take $\tilde{{\boldsymbol{v}}}={\boldsymbol{v}}_{1},\,\tilde{{\boldsymbol{u}}}={\boldsymbol{u}}_{1}$ to obtain from (3.12)
$\partial_{t}\rho_{1}+div_{x}(\rho_{1}({\boldsymbol{v}}+{\boldsymbol{v}}_{1}))=0.$} (3.16)
Furthermore, from (3.14) upon elimination of the constant $C$ using Lemma 3.1(iii)(iii), we have in $P_{1}$:
$\partial_{t}(\rho_{1}({\boldsymbol{v}}+{\boldsymbol{v}}_{1}))+div_{x}(\rho_{1}({\boldsymbol{u}}+{\boldsymbol{u}}_{1}))+{\nabla_{x}p(\rho_{1})+div_{x}\left[\rho_{1}(({\boldsymbol{v}}+{\boldsymbol{v}}_{1})\otimes({\boldsymbol{v}}+{\boldsymbol{v}}_{1})-({\boldsymbol{u}}+{\boldsymbol{u}}_{1})\right]}=0.$ (3.17)
The second and last terms in (3.17) cancel and we conclude that $(\rho_{1},\,{\boldsymbol{v}}+{\boldsymbol{v}}_{1})$ satisfy (3.1) and (3.2) in $P_{1}$ [^1] The derivation given here corrects the one given in our earlier paper [9].. An infinite number of weak solutions therefore has been obtained to the Riemann initial value problem with initial conditions (3.5). We emphasise that this set of weak solutions has been obtained for each fan sub-solution.
3.6 Parameterisation of admissible fan sub-solutions
Description of admissible fan sub-solutions is dealt with in [2, Section 4]. The conclusion we require is that the fan sub-solutions can be parameterised by $\rho_{1}$ and an additional parameter $\varepsilon_{2}\geq 0$; see [2, p. 1035]. Note that $\varepsilon_{2}$ has to satisfy the inequalities [2, (4.82),(4.83)] for admissibility.
We can express the constant $C$ of Lemma 3.1 in terms of these two parameters: the trace of the expression given in Lemma 3.1 $(iii)$ when $\Lambda=P_{1}$ yields
$C=|\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}}|^{2}\qquad\text{{in $P_{1}$.}}$ (3.18)
The constant $C$ in (3.18)) is then given by [2, p. 1043]:
$C=\beta^{2}(\rho_{1})+\varepsilon_{1}(\rho_{1})+\varepsilon_{2},$ (3.19)
where $\beta(\rho_{1})$ is the second component of ${\boldsymbol{v}}_{1}$ and is given by [2, (4.48), (4.53)] and $\varepsilon_{1}(\rho_{1})$ is given by [2, (4.49), (4.50)]. Both $\beta$ and $\varepsilon_{1}$ are smooth functions of $\rho_{1}$,
In the following, $\rho_{m}$ and $v_{m}$ refer, respectively, to the density and second component of the velocity for the intermediate state for the one-dimensional classical 2-shock solution to the Riemann initial value isentropic problem described in Section 3.2. We have (see [2, Lemma 4.5 and p. 1042]):
$\beta^{2}(\rho_{m})=v^{2}_{m},\qquad\varepsilon_{1}(\rho_{m})=0,$ (3.20)
and $\varepsilon_{2}>0$ on some interval
$I^{\ast}=[\rho^{\ast},\rho_{m}],$ (3.21)
where $\rho^{\ast}$ depends on the Riemann data. We also note that the 2-shock propagates with speeds $\nu_{\pm}(\rho_{m})$; see the discussion below (5.10) in [2].
3.7 Summary
Given the appropriate Riemann initial data, Chiodaroli and Kreml construct admissible fan sub-solutions parameterised by $\rho_{1}$ and $\varepsilon_{2}$. The domain of admissibility of fan sub-solutions is a nonempty domain $G$ in the $(\rho_{1},\varepsilon_{2})$ plane such that $G\subset(\rho_{+},\rho_{m})\times{\mathbb{R}}_{+}$. To each point in $G$ corresponds an uncountable set of convex integration (wild) solutions, which are physically indistinguishable: they all satisfy the same Rankine-Hugoniot conditions and have the same rate of energy dissipation.
Following [2] we denote by $\nu_{-}(\rho_{1}),\nu_{+}(\rho_{1})$ the shock speeds in the $x_{2}$ direction of all wild solutions corresponding to a particular admissible choice of $\rho_{1}$. Explicit expressions are presented in [2, (4.46),(4.47)].
The following sections apply the admissibility criterion LAAP0, formulated in Section 2, to the two-dimensional compressible isentropic Euler equations in the class $S$ consisting of the classical 2-shock solution, the convex integration solutions of [2] and the hybrid solutions of [14].
4 Justification for LAAP0
For some sub-solutions, the wild solutions constructed in [2] dissipate energy at $t=0^{+}$ more rapidly than the one-dimensional 2-shock solution. As a result, for some of their sub-solution data the 2-shock solution is not selected by the entropy rate criterion.
However, in the same Riemann initial value problem, LAAP prefers the 2-shock solution to the entropic wild solutions constructed in [2]. Nevertheless, Markfelder and Pellhammer [14] used Lemma 3.1 to construct solutions which are preferred by LAAP to the 2-shock solution. To this end, they restrict the two dimensional component of the fan sub-solution to a compact time interval. Consequently, for time $T_{0}$ greater than the upper limit of this compact time interval, a new problem may be constructed with piece-wise constant initial data determined by the fan sub-solution. The classical method of elementary waves is then used to extend locally in time the solution obtained in [2]. Hence Markfelder and Pellhammer [14] obtain hybrid solutions whose action for sufficiently large finite times is smaller than that of the 2-shock solution. It is worth remarking that the hybrid solution constructed by Markfelder and Pellhammer cannot be trivially extended to a global in time solution as the initial data at $t=T_{0}$ do not necessarily satisfy the small bounded variation hypothesis of Glimm’s global existence theorem [10].
These developments motivate a new admissibility criterion to replace the one formulated in [9]. The new feature is that the actions are defined with variable terminal time $t_{1}$ that enables the limit $t_{1}\rightarrow t_{0}^{+}$ to be taken.
5 Application of LAAP0
Recall that the Lagrangian $\mathcal{L}$ of fluid motion is the pointwise difference between the kinetic and potential energy and for the barotropic compressible fluid is given by
$\mathcal{L}=\frac{1}{2}\rho\left|{\boldsymbol{v}}\right|^{2}-\rho\varepsilon(\rho),$ (5.1)
in the notation of (3.4).
The corresponding time-dependent action $A(\rho,\,{\boldsymbol{v}})$ in a domain $\Lambda=\Omega\times[t_{o},\,t]$ is given by
$A(\rho,\,{\boldsymbol{v}})(t_{0},\,t)=\int^{t}_{t_{0}}\int_{\Omega}\mathcal{L}\,dx\,d\tau,\qquad t>t_{0}.$
The convex integration solution involves the constant $\rho_{1}$ introduced in (3.10) which we now restrict to a sufficiently small left interval $I$ contained in $I^{\ast}$ defined in (3.21). Put
[…]
and note that the 2-shock solution becomes identical to the convex integration solution of [2] in the region external to $\Lambda=P_{1}=[-L_{3},\,L_{3}]\times[\ell_{1},\,\ell_{2}]\times[0,\,t]$, where $2L_{3}$ is defined in Remark 3.1(ii)(ii), and
$\ell_{1}<\min{(\nu_{-}(\rho_{1}{)}:\rho_{1}\in I)}t\ {=\nu_{-}(\rho^{\ast})t},\qquad\ell_{2}>\max{(\nu_{+}(\rho_{1}{)}:\rho_{1}\in I)}t\ {=\nu_{+}(\rho^{\ast})t}.$ (5.3)
See Figure 1. The identities involving min, respectively max, here follow from the fact that $\nu_{-}$ is monotone increasing and $\nu_{+}$ is monotone decreasing on $I$, which in turn follows from equations (4.39)-(4.52) of [2].
Figure 1: Domains of integration
Set
$D(\rho_{1},\,t)=\int^{L_{3}}_{-L_{3}}\int^{t}_{0}\int^{\ell_{2}}_{\ell_{1}}L_{diff}(\rho_{1})\,dx_{2}d\tau dx_{1}.$
The function $D(\rho_{1},\,t)$ is continuous with respect to $\rho_{1}\in I$. Choose $\rho_{1}=\rho_{m}$ to obtain
$D(\rho_{m},t)=\int^{L_{3}}_{-L_{3}}\int^{t}_{0}\int^{\nu_{+}(\rho_{m})\tau}_{\nu_{-}(\rho_{m})\tau}L_{diff}(\rho_{m})\,dx_{2}d\tau dx_{1}.$
Observe that in the center wedge in Figure 1 equation (3.18) and (5.1) yield
$L_{diff}(\rho_{1})=\frac{1}{2}\rho_{m}v^{2}_{m}-\rho_{m}\varepsilon(\rho_{m})-\frac{1}{2}\rho_{1}C+\rho_{1}\varepsilon(\rho_{1}),$ (5.4)
where $C$ is given by (3.19). Hence using (3.19) and (3.20) we see that at $\rho_{1}=\rho_{m}$,
$L_{diff}(\rho_{m})=-\frac{1}{2}\rho_{m}\varepsilon_{2}<0.$
We conclude that $D(\rho_{m},t)<0$ for $t>0$. By continuity of $D(\rho_{1},t)$ we have $D(\rho_{1},t)<0$ for $t>0$ when $\rho_{1}$ belongs to a sufficiently small left neighbourhood of $\rho_{m}$. Further, because at $t=0$ the domain of integration shown in Figure 1, where the integrand is non-zero, collapses to a set of measure zero, we have
$\frac{\partial}{\partial t}D(\rho_{1},t)|_{t=0^{+}}=0.$ (5.5)
Finally, we note that
$\frac{\partial^{2}}{\partial t^{2}}D(\rho_{1},t)|_{t=0^{+}}<0$ (5.6)
in a sufficiently small left neighbourhood $I$ of $\rho_{m}$, because this inequality holds at $\rho_{1}=\rho_{m}$ and the second time derivative of $D(\rho_{1},t)$ is continuous in $\rho_{1}$.
We now state and prove the main results.
Consider the Riemann initial value problem for the two-dimensional isentropic compressible Euler system (3.1), (3.2) and (3.5). Let the set $S$ of solutions consist of the one-dimensional 2-shock solution, the entropic global convex integration weak solutions constructed by Chiodaroli and Kreml [2], and the hybrid solutions constructed by Markfelder and Pellhammer [14] with fan sub-solution data $\rho_{1}\in I^{\ast}$ and $\varepsilon_{2}>0$. Recall that for small $t>0$ the hybrid solutions coincide with those of [2]. We have
Theorem 5.1.
Given Riemann data satisfying (3.6), let $\varepsilon_{2}>0$ and assume that $\rho_{1}$ lies in a sufficiently small left neighbourhood $I$ of $\rho_{m}$. Then the 2-shock solution is sARAC-admissible, and therefore is the strictly LAAP0-admissible solution in $S$.
Proof.
The proof follows by combining the proof of [9, Thm. 5] with Definitions 2.2 and 2.3.
As just shown, for $\rho_{1}$ in the interval $I$ the action and its first time derivative are less than or equal and the second time derivative strictly less for the 2-shock solution than for any other solution in $S$, including those obtained in [14] since for small enough time these coincide with those of [2]. Hence, both sARAC and strict LAAP0 exclude the admissibility of the convex integration and hybrid solutions and establish the strict admissibility of the 2-shock solution in $S$. ∎
The next result is a global version of Theorem 5.1 corresponding to [9, Thm. 6]. It assumes the pressure law $p(\rho)=\rho^{\gamma}$ with $\gamma=2$ so that the internal energy is $\varepsilon=\rho$. Numerical evidence indicates that an analogous result holds for all $\gamma\in[1,\,3].$
Theorem 5.2.
Let $\gamma=2.$ For any Riemann data satisfying (3.6) and any $\varepsilon_{2}>0$ in (3.19), the classical 2-shock solution is sARACsARAC- and strictly LAAP0-admissible in $S$.
Proof.
The proof is a straightforward adaptation of [9, Thm. 6] and Theorem 5.1. ∎
Acknowledgements
The authors are grateful for support from the ICMS Research in Groups scheme. They also thank C. M. Dafermos, B. Gebhard and O. Kreml for remarks and the referees for their comments and corrections to an earlier version of this paper.